### v2

# Slope Consistency of Quasi-Maximum Likelihood

Abstract
Although QMLE is generally inconsistent, logistic regression relying on the binary choice model (BCM) with logistic errors is widely used, especially in machine learning contexts with many covariates and high-dimensional slope coefficients. This paper revisits the slope consistency of QMLE for BCMs. Ruud (1983) introduced a set of conditions under which QMLE may yield a constant multiple of the slope coefficient of BCMs asymptotically. However, he did not fully establish slope consistency of QMLE, which requires the existence of a positive multiple of slope coefficient identified as an interior maximizer of the population QMLE likelihood function over an appropriately restricted parameter space. We fill this gap by providing a formal proof of slope consistency under the same set of conditions for any binary choice model identified as in Manski (1975, 1985). Our result implies that logistic regression yields a consistent estimate for the slope coefficient of BCMs under suitable conditions.
JEL Classification: C13, C22
Keywords and phrases: binary choice models, quasi-maximum likelihood estimation, slope consistency, logistic regression
1 Introduction
Logistic regression is used extensively in machine learning to analyze binary outcomes associated with a large number of covariates and frequently adopted in practice as a quasi-maximum likelihood estimator (QMLE) for conventional binary choice models (BCMs). Logistic regression assumes the BCM with logistic errors, i.e., the logit model, although this is not explicitly recognized when used in machine learning contexts, which are generally astructural and not based on any fully specified model. The use of logistic regression or any other QMLE to estimate a BCM is generally invalid, yielding inconsistent estimates unless the error term in the BCM follows the same distribution as that employed in the QMLE. Furthermore, there are many other existing approaches to consistently estimate BCMs proposed by various authors. [^1] They include Powell, Stock, and Stoker (1989), Ichimura (1993), Klein and Spady (1993), Ahn, Ichimura, Powell, and Ruud (2018), Khan, Lan, Tamer, and Yao (2024), among others. Nevertheless, logistic regression is still used more widely in actual applications than any of these alternatives, perhaps due to its computational simplicity and the availability of software packages.
In this paper, we study consistency of QMLE for BCMs. Ruud (1983) provided conditions under which QMLE for BCMs may asymptotically yield a slope coefficient that is a constant multiple of the true slope coefficient. However, he did not fully establish the slope consistency of QMLE, which requires the existence of a positive multiple of the slope coefficient identified as an interior maximizer of the population QMLE likelihood function over an appropriately restricted parameter space. He showed neither that a solution to the first-order conditions for the maximizer of the likelihood function exists within the restricted parameter space, nor that the multiplicative constant is strictly positive. There may be no interior solution over the restricted parameter set. The multiplicative constant may also be zero or even negative, which could lead to incorrect conclusions of no effect or a reverse effect, respectively. We fill this gap by providing a formal proof of slope consistency under the same set of conditions for any binary choice model identified as in Horowitz (1992) or Manski (1975, 1985).
More specifically, we consider a binary outcome $Y$ given by the sign of $Y^{\ast}=V-U=\alpha_{0}+X^{\prime}\beta_{0}-U$, where $X$ is an $m$-dimensional covariate, and $V$ and $U$ are the scalar-valued systematic component and error term, respectively. For this BCM, we assume the index dependence given by $\mathcal{L}(U|X)=\mathcal{L}(U|V)$ with $\mathcal{L}$ denoting distribution and the linearity in expectation given by $\mathbb{E}(X|V)=aV+b$ with $a,b\in\mathbb{R}^{m}$. [^2] The index dependence is commonly imposed in the literature on consistent estimation of index models including all papers cited in Footnote 1. Ruud (1983) assumed the linearity in expectation, as well as the independence of $X$ and $U$, in which case the index dependence is also trivially satisfied. We establish slope consistency of QMLE with no other restrictive assumptions, except for necessary identification conditions and high-level technical conditions such as concavity and differentiability of the likelihood function which are commonly imposed to establish consistency of the maximum likelihood estimator (MLE). The linearity in expectation is certainly restrictive, but it holds, for instance, if $X$ has an elliptical distribution. Moreover, we may weight observations on $X$ appropriately so that the weighted observations can be regarded as a distribution satisfying the condition. See Ruud (1986) and Newey and Ruud (1994).
For binary outcomes generated by a BCM, QMLE is thus a viable option to use. In fact, logistic regression is used extensively in machine learning to analyze binary outcomes associated with a large number of covariates. This paper shows that it actually yields a consistent slope coefficient of the underlying binary choice model from which the binary outcomes are generated, provided that it satisfies the required index dependence and linearity in expectation conditions, in addition to other standard technical conditions. Needless to say, for a high-dimensional binary choice model, it is crucial to estimate the slope coefficient consistently. Once the slope coefficient is estimated consistently, it is relatively simple and straightforward to consistently estimate the intercept, which is one-dimensional.
The rest of the paper is organized as follows. In Section 2, we introduce our model and background theory with assumptions that are needed to ensure identification of the model and regularity in maximum likelihood estimation. Section 3 provides our main results on slope consistency of the QMLE with required assumptions on index dependence and linearity in expectation. Section 4 concludes the paper. Mathematical proofs are in the Appendix.
2 Preliminaries
Consider the binary choice model (BCM) given by
$Y=\operatorname*{sgn}\,(Y^{\ast})\quad\mbox{with}\quad Y^{\ast}=\alpha_{0}+X^{\prime}\beta_{0}-U,$ (1)
where $\operatorname*{sgn}$ is the sign function defined as $\operatorname*{sgn}\,(z)=\pm 1$ for $z\geq 0$ and $z<0$, respectively, $X$ is an $m$-dimensional vector of covariates, $\theta_{0}=(\alpha_{0},\beta_{0}^{\prime})^{\prime}$ is the true value of $(m+1)$-dimensional parameter vector $\theta=(\alpha,\beta^{\prime})^{\prime}$, and $U$ is the error term. [^3] We may define $Y=1\{Y^{\ast}\geq 0\}$ so that $Y$ takes value 1 or 0, following the usual convention for binary choice model. This will not change any of our subsequent discussions in the paper. In the paper, we define $Y$ to take value 1 or 1-1 to make our theory more directly applicable for the logistic regression used in machine learning. Throughout the paper, we let $\theta\in\Theta$, where $\Theta$ is a convex subset of $\mathbb{R}^{m+1}$ that is arbitrarily large. [^4] This is necessary because we identify $\theta_{0}$ only up to a scalar multiplication in the derivation of our main result. Our definition of $\Theta$ is essentially equivalent to set $\Theta=\mathbb{R}^{m+1}$. In the paper, however, we define $\Theta$ separately from $\mathbb{R}^{m+1}$ to make our subsequent assumptions more directly comparable to those existing in the literature.
Since we focus on the consistency of slope coefficient up to a positive scalar—referred to as slope consistency for simplicity—we impose conditions to ensure that $\theta_{0}$ is identified up to a positive scalar. Throughout the paper, we use $\mathcal{L}$ to denote the law or distribution.
Assumption 2.1.
$\operatorname*{med}\,(U|X)=0$ almost surely in $\mathcal{L}(X)$.
Assumption 2.2.
(a) $X_{m}$ has a nonzero coefficient, and the distribution of $X_{m}$ conditional on $X_{-m}$ has everywhere positive Lebesgue density almost surely in $\mathcal{L}(X_{-m})$, where $X_{-m}=(X_{1},\ldots,X_{m-1})^{\prime}$, (b) $0<\mathbb{P}\{Y=1|X\}<1$ almost surely in $\mathcal{L}(X)$, and (c) the support $\mathcal{X}$ of $\mathcal{L}(X)$ is not contained in any proper linear subspace of $\mathbb{R}^{m}$.
Assumptions 2.1 and 2.2 are imposed for the BCM in Horowitz (1992), modified from Manski (1975, 1985), which ensure that $\theta_{0}=(\alpha_{0},\beta_{0}^{\prime})^{\prime}$ is identified up to multiplication by a positive scalar with $\beta_{0}\neq 0$.
We consider the maximum likelihood estimation of a possibly misspecified model, referred to as quasi-maximum likelihood (QML) estimation, which assumes that $U$ is independent of $X$ and has distribution function $F$. The QML estimator $\hat{\theta}$ is defined as the maximum of
$Q_{n}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\Big{(}1\{Y_{i}=1\}\log F(\alpha+X_{i}^{\prime}\beta)+1\{Y_{i}=-1\}\log\big{(}1-F(\alpha+X_{i}^{\prime}\beta)\big{)}\Big{)}$ (2)
over a parameter set $\Theta$. We define
$Q(\theta)=\mathbb{E}\Big{(}1\{Y=1\}\log F(\alpha+X^{\prime}\beta)+1\{Y=-1\}\log\big{(}1-F(\alpha+X^{\prime}\beta)\big{)}\Big{)}$ (3)
Assumption 2.3.
(a) $Q$ has a maximum $\theta_{\ast}$ which is an interior point of $\Theta$, (b) $\mathbb{E}|\log F(\alpha+X^{\prime}\beta)|,\mathbb{E}\big{|}\log\big{(}1-F(\alpha+X^{\prime}\beta)\big{)}\big{|}<\infty$ for any $\theta\in\Theta$, and (c) $F(\cdot)\in(0,1)$ on $\mathbb{R}$, and $\log F(\cdot),\log(1-F(\cdot))$ are strictly concave.
Assumption 2.3 is standard. Part (a) assumes the existence of a pseudo-true value given by QMLE, which is commonly imposed in the existing work of QMLE, see, e.g., White (1982), Gourieroux, Monfort, and Trognon (1984). Part (b) ensures that $Q(\theta)$ is well-defined and $\hat{Q}(\theta)\to_{p}Q(\theta)$ for any $\theta\in\Theta$. Finally, Part (c) guarantees that $\theta_{\ast}$ is unique.
Lemma 2.1.
Let Assumption 2.3 hold. Then $\hat{\theta}\to_{p}\theta_{\ast}$.
Lemma 2.1 follows immediately from Theorem 2.7 in Newey and McFadden (1994), since all conditions required in their theorem are trivially satisfied under our Assumption 2.3.
Lemma 2.1 shows that $\hat{\theta}$ has a probability limit $\theta_{\ast}$. Similar results are provided by White (1982) and Gourieroux, Monfort, and Trognon (1984). Unlike these works, we assume neither the compactness of $\Theta$ nor the existence of an envelope function for $|\log F(\alpha+x^{\prime}\beta)|$ and $|\log(1-F(\alpha+x^{\prime}\beta))|$ with finite moment. Instead, we assume convexity of $\Theta$ and impose concavity assumptions on $\log F$ and $\log(1-F)$.
Assumption 2.4.
The derivative $f$ of $F$ is well defined and continuously differentiable on $\mathbb{R}$.
Given Assumption 2.3, Assumption 2.4 implies that $\theta_{\ast}$ is uniquely defined as the solution to the first order condition (FOC) given by the derivative of $Q$. Both Part (c) of Assumption 2.3 and Assumption 2.4 are satisfied for normal and logistic distribution functions, which are most commonly used in practice.
Subsequently, we let Assumption 2.4, as well as Assumption 2.3, hold. Under these assumptions, $Q$ is strictly concave and twice continuously differentiable. Therefore, if we define
$\dot{\ell}_{+}(z)=\frac{f(z)}{F(z)}\quad\text{and}\quad\dot{\ell}_{-}(z)=\frac{f(z)}{1-F(z)},$ (4)
the probability limit $\theta_{\ast}=(\alpha_{\ast},\beta_{\ast}^{\prime})^{\prime}$ of $\hat{\theta}$ is defined as the solution to the first-order condition $\dot{Q}(\theta)=0$, where $\dot{Q}$ is the first-order derivative of $Q$ given by
$\dot{Q}(\theta)=\mathbb{E}\left(1\{Y=1\}\dot{\ell}_{+}(\alpha+X^{\prime}\beta)-1\{Y=-1\}\dot{\ell}_{-}(\alpha+X^{\prime}\beta)\right)\begin{pmatrix}1\\ X\end{pmatrix}$. (5)
[…]
3 Slope Consistency
Let
$V=\alpha_{0}+X^{\prime}\beta_{0}$
in what follows. We introduce two important assumptions that are crucial for the slope consistency of QMLE.
Assumption 3.1.
$\mathcal{L}(U|X)=\mathcal{L}(U|V)$.
Assumption 3.1 requires that the error distribution depends on $X$ only through the index $V$. Such an index dependence condition for the error distribution is often used in the literature. It’s used in, e.g., Klein and Spady (1993) for the BCM, and Ichimura (1993) for a class of single index models, among many others. This condition is satisfied trivially if $U$ is independent of $X$, as is often assumed for the BCM.
Assumption 3.2.
$\mathbb{E}(X|V)=aV+b$ for some $a,b\in\mathbb{R}^{m}$.
Assumption 3.2 requires that the conditional mean of $X$ given $V$ is a linear function of $V$, which will be referred to simply as the linearity in expectation condition. This condition is certainly restrictive, but it holds, for instance, if $X$ has an elliptical distribution. We may also weight observations on $X$ appropriately so that the distribution of the weighted observations can be regarded as satisfying this condition. See, e.g., Ruud (1986) and Newey and Ruud (1994). In place of original observations $\{x_{i}\}_{i=1}^{n}$ on $X$, they use the observations weighted by $w_{i}=\sigma(x_{i})/\hat{\tau}(x_{i})$ for $i=1,\ldots,n$, where $\sigma$ is the standard multivariate normal density satisfying Assumption 3.2, and $\hat{\tau}$ is a local kernel density estimator of $X$. Note that the weighted observations $\{w_{i}x_{i}\}_{i=1}^{n}$ may be regarded approximately as the observations drawn from $X$ having density $\sigma$. This was proposed by Ruud (1986) for QMLE, and subsequently used by Newey and Ruud (1994) for weighted least squares. [^5] Ruud (1986) studies a class of limited dependent variable models given by $Y=\pi(\alpha_{0}+X^{\prime}\beta_{0}-U)$ with known $\pi$, and Newey and Ruud (1994) consider the semiparametric single index model given by $\mathbb{E}(Y|X)=\pi(X^{\prime}\beta_{0})$ with unknown $\pi$.
For slope consistency of QMLE, we consider the QMLE over a restricted set of parameters $\theta=(\alpha,\beta^{\prime})^{\prime}$ specified as
$\begin{pmatrix}\alpha\\ \beta\end{pmatrix}=c\begin{pmatrix}\alpha_{0}\\ \beta_{0}\end{pmatrix}+\begin{pmatrix}r\\ 0\end{pmatrix}$ (6)
with some $c,r\in\mathbb{R}$. The resulting estimator will be referred to simply as the restricted QMLE. Note that the restricted QMLE has only two parameters, while the unrestricted QMLE has $(m+1)$-parameters. Let $\dot{Q}(c,r)$ be defined as $\dot{Q}(\theta)$ in (5) with the $(m+1)$-dimensional parameter $\theta=(\alpha,\beta^{\prime})^{\prime}$ being restricted to the two dimensional parameter $(c,r)$ introduced in (6). Moreover, we let
$\Pi(v)=\mathbb{P}\big{\{}Y=1\big{|}V=v\big{\}}=\mathbb{P}\big{\{}U\leq V\big{|}V=v\big{\}}$
for $v\in\mathbb{R}$. Under Assumption 2.1, we have $\Pi(v)\geq 1/2$ if $v\geq 0$, and $\Pi(v)<1/2$ if $v<0$.
It follows that
$\dot{Q}(c,r)$ $=\mathbb{E}\left(1\{U\leq V\}\dot{\ell}_{+}(cV+r)-1\{U>V\}\dot{\ell}_{-}(cV+r)\right)\begin{pmatrix}1\\ X\end{pmatrix}$
$=\mathbb{E}\left(\Pi(V)\dot{\ell}_{+}(cV+r)-(1-\Pi(V))\dot{\ell}_{-}(cV+r)\right)\begin{pmatrix}1\\ \mathbb{E}(X|V)\end{pmatrix}$
$=\mathbb{E}\left(\Pi(V)\dot{\ell}_{+}(cV+r)-(1-\Pi(V))\dot{\ell}_{-}(cV+r)\right)\begin{pmatrix}1\\ aV+b\end{pmatrix}$, (7)
where we use the law of iterated expectation by taking conditional expectation on $V$ before we take expectation $\mathbb{E}$ to define $\dot{Q}(c,r)$, and apply the linearity in expectation for $\mathbb{E}(X|V)$ in Assumption 3.2. Consequently, we may deduce from (7) that the FOC for $(c,r)$ given by the system of $(m+1)$-equations $\dot{Q}(c,r)=0$ effectively reduces to the following system of two equations
$\dot{Q}_{\bullet}(c,r)$ $=\mathbb{E}\left(1\{U\leq V\}\dot{\ell}_{+}(cV+r)-1\{U>V\}\dot{\ell}_{-}(cV+r)\right)\begin{pmatrix}1\\ V\end{pmatrix}$
$=\mathbb{E}\left(\Pi(V)\dot{\ell}_{+}(cV+r)-(1-\Pi(V))\dot{\ell}_{-}(cV+r)\right)\begin{pmatrix}1\\ V\end{pmatrix}=0$ (8)
with two unknowns $(c,r)$.
As discussed, Assumption 2.3 ensures that the likelihood function for the unrestricted QMLE has a unique interior maximizer asymptotically. This, however, does not mean that the likelihood function for the restricted QMLE also has a unique interior maximizer asymptotically. Therefore, we cannot automatically assume that the solution to the FOC in (3) exists. In fact, the slope consistency of QMLE requires that the FOC has a solution $(c_{\ast},r_{\ast})$ with $c_{\ast}>0$. Note that any solution $(c_{\ast},r_{\ast})$ to the FOC is a maximizer of the unrestricted QMLE, and therefore, it is a unique solution under Assumptions 3.1 and 3.2. In fact, Ruud (1983) shows neither that the FOC has a solution $(c_{\ast},r_{\ast})$, nor that $c_{\ast}>0$. He simply assumes that the likelihood function for his restricted QMLE, which is essentially identical to ours, has the maximizer given by a solution to the FOC. Li and Duan (1989) consider a more general class of models and introduce an additional high-level assumption which is sufficient to ensure for binary choice model that the FOC has a solution $c_{\ast}\in\mathbb{R}$, but not necessarily with $c_{\ast}>0$. See also Powell (1994) for related discussions.
Lemma 3.1.
Let Assumptions 2.1, 2.2, 2.3 and 2.4 hold. Then $\dot{Q}_{\bullet}(c,r)=0$ has a solution $(c_{\ast},r_{\ast})$ such that $c_{\ast}>0$ and $r_{\ast}\in\mathbb{R}$.
Theorem 3.2.
Let Assumptions 2.1, 2.2, 2.3, 2.4, 3.1 and 3.2 hold. Then $\dot{Q}_{\bullet}(c,r)=0$ has a unique solution $(c_{\ast},r_{\ast})$ for $c_{\ast}>0$ and $r_{\ast}\in\mathbb{R}$. Moreover, if we let $\theta_{\ast}=(\alpha_{\ast},\beta_{\ast}^{\prime})^{\prime}$ with $\alpha_{\ast}=c_{\ast}\alpha_{0}+r_{\ast}$ and $\beta_{\ast}=c_{\ast}\beta_{0}$, then $\hat{\theta}\to_{p}\theta_{\ast}$ as $n\to\infty$.
Theorem 3.2, which follows immediately from Lemma 3.1 under additional conditions in Assumptions 3.1 and 3.2, establishes slope consistency of QMLE for binary choice models.
4 Concluding Remarks
We establish the slope consistency of the QMLE for binary choice models (BCMs) under suitable conditions, which provides some theoretical justification for the widespread use of logistic regression in machine learning for analyzing binary outcomes as well as the popularity of logit and probit models in applied work. We focus on consistency of the slope parameters of BCMs because they are of primary interest. Empirical analyses often rely on the relative magnitudes of slope coefficients to assess the role of covariates in latent utilities. Moreover, when the covariate is high-dimensional, as is the case in many empirical applications nowadays, slope consistency is particularly essential, since the remaining one-dimensional intercept parameter can be estimated easily once an estimate of the slope parameter is obtained.
Our conditions for slope consistency are not as restrictive as they may appear. The index dependence condition is standard in the literature and weaker than assuming independent errors. More importantly, the linearity-in-expectation condition holds if covariates jointly have an elliptical distribution, and even when this is not the case, we may assign appropriate weights to the observations on covariates so that they may be regarded as being drawn from such a distribution. See, e.g., Ruud (1986), Newey and Ruud (1994). Thus, slope consistency of QMLE is not only of theoretical interest but also practically attainable under a broader range of data-generating processes than might be expected.
Appendix: Mathematical Proofs
A.1 Proof of Lemma 3.1
The proof of Lemma 3.1 will be done in three steps. We let
$\phi(c,r)$ $=\mathbb{E}\Big{(}1\{U\leq V\}\dot{\ell}_{+}(cV+r)-1\{U>V\}\dot{\ell}_{-}(cV+r)\Big{)}$
$=\mathbb{E}\Big{(}\Pi(V)\dot{\ell}_{+}(cV+r)-\big{(}1-\Pi(V)\big{)}\dot{\ell}_{-}(cV+r)\Big{)}$
and
$\psi(c,r)$ $=\mathbb{E}\Big{(}1\{U\leq V\}\dot{\ell}_{+}(cV+r)-1\{U>V\}\dot{\ell}_{-}(cV+r)\Big{)}V$
$=\mathbb{E}\Big{(}\Pi(V)\dot{\ell}_{+}(cV+r)-\big{(}1-\Pi(V)\big{)}\dot{\ell}_{-}(cV+r)\Big{)}V,$
and show that the two equations given by $\phi(c,r)=0$ and $\psi(c,r)=0$ have a solution $(c_{\ast},r_{\ast})$ with $c_{\ast}>0$ and $r_{\ast}\in\mathbb{R}$. In the first step, we show that there exists a unique $r(c)\in\mathbb{R}$ such that $\phi(c,r(c))=0$ for any $c\geq 0$. Then we let
$\psi(c)=\psi\big{(}c,r(c)\big{)},$
and show that $\psi(0)>0$ and $\psi(c)<0$ for some $c>0$ in the second and third steps, respectively. The first and second steps are straightforward, while the third step is more involved. Throughout the proofs, we repeatedly use the fact that $\dot{\ell}_{+}(z)\big{(}\dot{\ell}_{-}(z)\big{)}$ is continuous, and strictly decreasing (increasing) to zero (from zero) on $\mathbb{R}$ as $z\to\infty$ with intersection at $z=0$ with their value $2f(0)$.
Step 1:
In this step, we fix $c\geq 0$ arbitrarily, and show that there exists a unique $r(c)\in\mathbb{R}$ such that $\phi(c,r(c))=0$. Note that
$\lim_{r\to-\infty}\phi(c,r)\in(0,\infty]\quad\mbox{and}\quad\lim_{r\to\infty}\phi(c,r)\in[-\infty,0),$
and that $\phi(c,r)$ is continuous and strictly decreasing in $r\in\mathbb{R}$ for any given $c\geq 0$. Therefore, for any $c\geq 0$, there exists a unique $r(c)\in\mathbb{R}$ such that $\phi(c,r(c))=0$ by the intermediate value theorem. In particular, $r(0)$ is defined as
$\phi(0,r(0))=\dot{\ell}_{+}\big{(}r(0)\big{)}\mathbb{P}\{U\leq V\}-\dot{\ell}_{-}\big{(}r(0)\big{)}\mathbb{P}\{U>V\}=0,$ (9)
which will be used subsequently.
Step 2:
In this step, we show that $\psi(0)>0$. This follows immediately, since
$\psi(0)$ $=\dot{\ell}_{+}\big{(}r(0)\big{)}\mathbb{E}\big{(}V1\{U\leq V\}\big{)}-\dot{\ell}_{-}\big{(}r(0)\big{)}\mathbb{E}\big{(}V1\{U>V\}\big{)}$
$=\dot{\ell}_{+}\big{(}r(0)\big{)}\mathbb{E}\big{(}V\big{|}U\leq V\big{)}\mathbb{P}\{U\leq V\}-\dot{\ell}_{-}\big{(}r(0)\big{)}\mathbb{E}\big{(}V\big{|}U>V\big{)}\mathbb{P}\{U>V\}$
$=\dot{\ell}_{+}\big{(}r(0)\big{)}\Big{(}\mathbb{E}\big{(}V\big{|}U\leq V\big{)}-\mathbb{E}\big{(}V\big{|}U>V\big{)}\Big{)}\mathbb{P}\{U\leq V\}>0$ (10)
[…]
Step 3:
In this step, we show that
$\psi(c)<0$ (11)
for some $c>0$. Here we define $\psi(c)$ equivalently as
$\psi(c)$ $=\mathbb{E}\Big{(}\Pi(V)\dot{\ell}_{+}(cV+r(c))-\big{(}1-\Pi(V)\big{)}\dot{\ell}_{-}(cV+r(c))\Big{)}\big{(}cV+r(c)\big{)}$
$=\mathbb{E}\Big{(}\Pi(V)\dot{\ell}_{+}\big{(}c(V+s(c))\big{)}-\big{(}1-\Pi(V)\big{)}\dot{\ell}_{-}\big{(}c(V+s(c))\big{)}\Big{)}\big{(}c(V+s(c))\big{)}$
with $s(c)=r(c)/c$ for $c>0$. Under this definition, we may easily see why (11) should hold. For any small $\delta>0$, there exists a large enough $c>0$ such that
$\Pi(v)\dot{\ell}_{+}\big{(}c(v+s(c))\big{)}-\big{(}1-\Pi(v)\big{)}\dot{\ell}_{-}\big{(}c(v+s(c))\big{)}$
$\quad\approx\begin{cases}-\big{(}1-\Pi(v)\big{)}\dot{\ell}_{-}\big{(}c(v+s(c))\big{)}<0&\text{when}\ v+s(c)>\delta\\ \Pi(v)\dot{\ell}_{+}\big{(}c(v+s(c))\big{)}>0&\text{when}\ v+s(c)<-\delta\end{cases},$
from which it follows that
$\Big{(}\Pi(v)\dot{\ell}_{+}\big{(}c(v+s(c))\big{)}-\big{(}1-\Pi(v)\big{)}\dot{\ell}_{-}\big{(}c(v+s(c))\big{)}\Big{)}(v+s(c))<0$
for all $v\in\mathbb{R}$.
In the rest of the proof, we will show it more rigorously. Let $\{c_{n}\}$ be a sequence of numbers such that $c_{n}>0$ for all $n$ and $c_{n}\to\infty$ as $n\to\infty$, and define $s_{n}=s(c_{n})$ for all $n$. Then it suffices to show that
$\liminf_{n\to\infty}\mathbb{E}\Big{(}\Pi(V)\dot{\ell}_{+}(c_{n}(V+s_{n}))-\big{(}1-\Pi(V)\big{)}\dot{\ell}_{-}(c_{n}(V+s_{n}))\Big{)}(V+s_{n})<0$ (12)
as $n\to\infty$, to establish (11). We may assume without loss of generality, by taking a subsequence if necessary, that the sequence $\{s_{n}\}$ is monotonic. Since $\{s_{n}\}$ is monotonic by construction, we have either (i) $s_{n}\to\pm\infty$ or (ii) $s_{n}\to\bar{s}\in\mathbb{R}$ as $n\to\infty$, depending upon whether $\{s_{n}\}$ is unbounded or bounded.
Since $f$ is a density function, we have
$f(z)=o\left(\frac{1}{|z|}\right)$
as $z\to\pm\infty$, from which it follows that
$\lim_{z\to\infty}\frac{\dot{\ell}_{+}(z)}{1/|z|}=\lim_{z\to\infty}\frac{|z|f(z)}{F(z)}=0,\quad\lim_{z\to-\infty}\frac{\dot{\ell}_{-}(-z)}{1/|z|}=\lim_{z\to-\infty}\frac{|z|f(z)}{1-F(z)}=0,$
and that
$\sup_{z\geq 0}\frac{\dot{\ell}_{+}(z)}{1/|z|}<\infty,\quad\sup_{z\leq 0}\frac{\dot{\ell}_{-}(z)}{1/|z|}<\infty.$
For any $\delta>0$ small, we therefore have
$\Big{|}\mathbb{E}1\{V+s_{n}>\delta\}\Pi(V)\dot{\ell}_{+}\big{(}c_{n}(V+s_{n})\big{)}c_{n}(V+s_{n})\Big{|}$ $\leq\sup_{z\geq 0}\frac{\dot{\ell}_{+}(z)}{1/|z|}<\infty$
$\Big{|}\mathbb{E}1\{V+s_{n}<-\delta\}\big{(}1-\Pi(V)\big{)}\dot{\ell}_{-}\big{(}c_{n}(V+s_{n})\big{)}c_{n}(V+s_{n})\Big{|}$ $\leq\sup_{z\leq 0}\frac{\dot{\ell}_{-}(z)}{1/|z|}<\infty,$
from which we may easily deduce that
$\mathbb{E}1\{V+s_{n}>\delta\}\Pi(V)\dot{\ell}_{+}\big{(}c_{n}(V+s_{n})\big{)}(V+s_{n})$ $\to 0$
$\mathbb{E}1\{V+s_{n}<-\delta\}\big{(}1-\Pi(V)\big{)}\dot{\ell}_{-}\big{(}c_{n}(V+s_{n})\big{)}(V+s_{n})$ $\to 0$
as $n\to\infty$.
It follows that
$\mathbb{E}1\{V+s_{n}>\delta\}\Big{(}\Pi(V)\dot{\ell}_{+}\big{(}c_{n}(V+s_{n})\big{)}-\big{(}1-\Pi(V)\big{)}\dot{\ell}_{-}\big{(}c_{n}(V+s_{n})\big{)}\Big{)}(V+s_{n})$
$=o(1)-\mathbb{E}1\{V+s_{n}>\delta\}\big{(}1-\Pi(V)\big{)}\dot{\ell}_{-}\big{(}c_{n}(V+s_{n})\big{)}(V+s_{n})$
$<o(1)-\mathbb{E}1\{V+s_{n}>\delta\}\big{(}1-\Pi(V)\big{)}\dot{\ell}_{-}(c_{n}\delta)\delta$ (13)
and
$\mathbb{E}1\{V+s_{n}<-\delta\}\Big{(}\Pi(V)\dot{\ell}_{+}\big{(}c_{n}(V+s_{n})\big{)}-\big{(}1-\Pi(V)\big{)}\dot{\ell}_{-}\big{(}c_{n}(V+s_{n})\big{)}\Big{)}(V+s_{n})$
$=o(1)+\mathbb{E}1\{V+s_{n}<-\delta\}\Pi(V)\dot{\ell}_{+}\big{(}c_{n}(V+s_{n})\big{)}(V+s_{n})$
$<o(1)-\mathbb{E}1\{V+s_{n}<-\delta\}\Pi(V)\dot{\ell}_{+}(-c_{n}\delta)\delta$ (14)
for all large $n$. Furthermore, we have
$\Big{|}\mathbb{E}1\{-\delta\!<\!V\!+\!s_{n}\!<\!0\}\big{(}1\!-\!\Pi(V)\big{)}\dot{\ell}_{-}\big{(}c_{n}(V\!+\!s_{n})\big{)}(V\!+\!s_{n})\Big{|}$ $\leq\mathbb{P}\{-\delta\!<\!V\!+\!s_{n}\!<\!0\}\dot{\ell}_{-}(0)\delta$
$\Big{|}\mathbb{E}1\{0\!<\!V\!+\!s_{n}\!<\!\delta\}\Pi(V)\dot{\ell}_{+}\big{(}c_{n}(V\!+\!s_{n})\big{)}(V\!+\!s_{n})\Big{|}$ $\leq\mathbb{P}\{0\!<\!V\!+\!s_{n}\!<\!\delta\}\dot{\ell}_{+}(0)\delta,$
from which we may deduce that
$\mathbb{E}1\{-\delta<V\!+\!s_{n}<\delta\}\Big{(}\Pi(V)\dot{\ell}_{+}\big{(}c_{n}(V\!+\!s_{n})\big{)}-\big{(}1-\Pi(V)\big{)}\dot{\ell}_{-}\big{(}c_{n}(V\!+\!s_{n})\big{)}\Big{)}(V\!+\!s_{n})$
$\leq\Big{(}\dot{\ell}_{-}(0)\mathbb{P}\{-\delta<V\!+\!s_{n}<0\}+\dot{\ell}_{+}(0)\mathbb{P}\{0<V\!+\!s_{n}<\delta\}\Big{)}\delta$
$\quad+\mathbb{E}1\{-\delta<V\!+\!s_{n}<0\}\Pi(V)\dot{\ell}_{+}\big{(}c_{n}(V\!+\!s_{n})\big{)}(V\!+\!s_{n})$
$\quad-\mathbb{E}1\{0<V\!+\!s_{n}<\delta\}\big{(}1-\Pi(V)\big{)}\dot{\ell}_{-}\big{(}c_{n}(V\!+\!s_{n})\big{)}(V\!+\!s_{n})$
$<\Big{(}\dot{\ell}_{-}(0)\mathbb{P}\{-\delta<V\!+\!s_{n}<0\}+\dot{\ell}_{+}(0)\mathbb{P}\{0<V\!+\!s_{n}<\delta\}\Big{)}\delta$ (15)
for all large $n$. Consequently, it follows from (A.1), (A.1) and (A.1) that
$\mathbb{E}\Big{(}\Pi(V)\dot{\ell}_{+}\big{(}c_{n}(V\!+\!s_{n})\big{)}-\big{(}1-\Pi(V)\big{)}\dot{\ell}_{-}\big{(}c_{n}(V\!+\!s_{n})\big{)}\Big{)}(V\!+\!s_{n})$
$<o(1)-\dot{\ell}_{-}(c_{n}\delta)\delta\mathbb{E}1\{V\!+\!s_{n}>\delta\}\big{(}1-\Pi(V)\big{)}-\dot{\ell}_{+}(-c_{n}\delta)\delta\mathbb{E}1\{V\!+\!s_{n}<-\delta\}\Pi(V)$
$\quad+\bigg{(}\dot{\ell}_{-}(0)\mathbb{P}\{-\delta<V\!+\!s_{n}<0\}+\dot{\ell}_{+}(0)\mathbb{P}\{0<V\!+\!s_{n}<\delta\}\Big{)}\delta$
$<o(1)-\delta\Big{(}\dot{\ell}_{-}(0)\mathbb{E}1\{V\!+\!s_{n}>\delta\}\big{(}1-\Pi(V)\big{)}+\dot{\ell}_{+}(0)\mathbb{E}1\{V\!+\!s_{n}<-\delta\}\Pi(V)$
$\qquad\qquad\qquad-\big{(}\dot{\ell}_{-}(0)+\dot{\ell}_{+}(0)\big{)}\mathbb{P}\{-\delta<V\!+\!s_{n}<\delta\}\bigg{)}$ (16)
for all large $n$.
If $s_{n}\to\pm\infty$, then $\mathbb{P}\{-\delta<V\!+\!s_{n}<\delta\}\to 0$. In this case, (A.1) yields
$\liminf_{n\to\infty}\mathbb{E}\Big{(}\Pi(V)\dot{\ell}_{+}\big{(}c_{n}(V\!+\!s_{n})\big{)}-\big{(}1-\Pi(V)\big{)}\dot{\ell}_{-}\big{(}c_{n}(V\!+\!s_{n})\big{)}\Big{)}(V\!+\!s_{n})$
$\leq\liminf_{n\to\infty}\bigg{[}-\delta\Big{(}\dot{\ell}_{-}(0)\mathbb{E}1\{V\!+\!s_{n}>\delta\}\big{(}1-\Pi(V)\big{)}+\dot{\ell}_{+}(0)\mathbb{E}1\{V\!+\!s_{n}<-\delta\}\Pi(V)\Big{)}\bigg{]}<0,$
since $\dot{\ell}_{+}(0),\dot{\ell}_{-}(0)>0$ and $\mathbb{E}\Pi(V)\in(0,1)$, and therefore, (12) holds. On the other hand, in case $s_{n}\to\bar{s}\in\mathbb{R}$, (A.1) yields
$\liminf_{n\to\infty}\mathbb{E}\Big{(}\Pi(V)\dot{\ell}_{+}\big{(}c_{n}(V\!+\!s_{n})\big{)}-\big{(}1-\Pi(V)\big{)}\dot{\ell}_{-}\big{(}c_{n}(V\!+\!s_{n})\big{)}\Big{)}(V\!+\!s_{n})$
$\leq-\delta\Big{(}\dot{\ell}_{-}(0)\mathbb{E}1\{V\!+\!\bar{s}>\delta\}\big{(}1-\Pi(V)\big{)}+\dot{\ell}_{+}(0)\mathbb{E}1\{V\!+\!\bar{s}<-\delta\}\Pi(V)$
$\qquad\qquad-\big{(}\dot{\ell}_{-}(0)+\dot{\ell}_{+}(0)\big{)}\mathbb{P}\{|V\!+\!\bar{s}|<\delta\}\Big{)}<0$
for all $\delta>0$ small enough, since $\mathbb{P}\{|V+\bar{s}|<\delta\}\to 0$ as $\delta\to 0$ due to Assumption 2.2 which implies that $V$ has a positive Lebesgue density on $\mathbb{R}$. Therefore, it follows that (12) holds also in this case.
With Steps 1, 2, and 3 established, we are now ready to prove Lemma 3.1. Since $\phi(c,r)$ is continuous in $(c,r)$, we may readily deduce that $r(c)$ is continuous in $c$. Likewise, $\psi(c,r)$ is continuous in $(r,c)$, from which it follows that $\psi(c)$ is continuous in $c>0$. Therefore, Step 2 and Step 3 imply that $\psi(c)=0$ has a solution $c_{\ast}>0$ by the intermediate value theorem. Once $c_{\ast}>0$ is obtained, we may let $r_{\ast}\in\mathbb{R}$ be given by $r_{\ast}=r(c_{\ast})$, where $r(c)$ is defined in Step 1. It is clear that $(c_{\ast},r_{\ast})$ is a solution to $\dot{Q}_{\bullet}(c,r)=0$, as was to be shown. This completes the proof.
A.2 Proof of Theorem 3.2
Since $(c_{\ast},r_{\ast})$ is a solution to $\dot{Q}_{\bullet}(c,r)=0$, it is also a solution to $\dot{Q}(c,r)=0$ under Assumptions 3.1 and 3.2. The identifying conditions in Assumptions 2.1 and 2.2 ensure that $(c_{\ast},r_{\ast})$ define $\theta_{\ast}=(\alpha_{\ast},\beta_{\ast}^{\prime})^{\prime}$ uniquely. However, due to Assumptions 2.3 and 2.4, $\theta_{\ast}$ defined by $(c_{\ast},r_{\ast})$ is uniquely given as the probability limit of QMLE. This implies that $(c_{\ast},r_{\ast})$ is unique. The proof is therefore complete.